Avogadro's number is 6.02 x 1023 (6.02E+23 in e-notation). How many molecules are there in 0.5 moles of glucose which has a molecular weight of 180 grams per mole? Express your answer in e-notation.
Answer #1

1 mole contains6.02*10^23 molecules

number of molecules= number of moles* Avogadro number

=0.5 * 6.02*10^23

= 3.01*10^23

so number of molecules= 3.01*10^23=3.01E+23